# Topology Sculptor, Shape Refiner: Discrete Diffusion Model for High-Fidelity 3D Meshes Generation

### v1

@@ Abstract
In this paper, we introduce Topology Sculptor, Shape Refiner (TSSR), a novel method for generating high-quality, artist-style 3D meshes based on Discrete Diffusion Models (DDMs). Our primary motivation for TSSR is to achieve highly accurate token prediction while enabling parallel generation, a significant advantage over sequential autoregressive methods. By allowing TSSR to ”see” all mesh tokens concurrently, we unlock a new level of efficiency and control. We leverage this parallel generation capability through three key innovations: 1) Decoupled Training and Hybrid Inference, which distinctly separates the DDM-based generation into a topology sculpting stage and a subsequent shape refinement stage. This strategic decoupling enables TSSR to effectively capture both intricate local topology and overarching global shape. 2) An Improved Hourglass Architecture, featuring bidirectional attention enriched by face-vertex-sequence level Rotational Positional Embeddings (RoPE), thereby capturing richer contextual information across the mesh structure. 3) A novel Connection Loss, which acts as a topological constraint to further enhance the realism and fidelity of the generated meshes. Extensive experiments on complex datasets demonstrate that TSSR generates high-quality 3D artist-style meshes, capable of achieving up to 10,000 faces at a remarkable spatial resolution of $1024^{3}$. The code will be released at: https://github.com/psky1111/Tencent-TSSR.
@@ 1 Introduction
Mainstream 3D mesh generation methods can be broadly categorized into two approaches: 1) Shape-then-mesh generation and 2) direct mesh generation. Shape-then-mesh generations [28] start with representing a shape via continuous volumetric fields, such as Neural Radiance Fields (NeRF) [17] and Signed Distance Fields (SDF) [10]. Then, a diffusion model is used to generate such a representation and convert it to real meshes via Marching Cubes [15]. Benefits from parallel token generation in diffusion models, but shape-then-mesh generation is efficient but suffers from inevitable denegation after the conversion process. Direct mesh generation methods aim to predict mesh vertices and faces directly. Mainstream methods, such as Meshtron [8] and DeepMesh [27], employ the Auto-Regressive (AR) approach, which first quantizes the mesh coordinates into a discrete space (V discrete bins for each coordinate, where V is also referred to as the spatial resolution in this context). Then it implements the specific tokenizer method to generate the mesh token-by-token. Benefits from the discrete space, ARs can generate high-quality meshes, but struggle with efficiency.
@@
In this paper, we propose the Topology Sculptor, Shape Refiner (TSSR), a novel paradigm explicitly designed for direct mesh generation built upon DDMs. TSSR addresses the challenge of maintaining accurate token prediction in parallel generation for meshes. TSSR leverages DDMs’ inherent ability for parallel token generation, which enables the model to view all tokens simultaneously and utilize bidirectional attention to capture comprehensive global shape and intricate topological information. TSSR realizes accurate parallel mesh generation through three integrated core contributions:
1) A unique decoupling training and hybrid inference paradigm that strategically divides mesh generation into specialized sub-tasks for shape (shape refiner) and topology (topology sculptor). This strategy acknowledges that while parallel generation yields rich, comprehensive information (encompassing global shape and intricate topology), effectively processing this intertwined complexity for accurate prediction is challenging. Decoupling training enables the separation of these aspects into distinct sub-tasks, thereby reducing the learning burden on the model and allowing each sub-task to focus predominantly on either shape or topology. Then, the hybrid inference process unifies these sub-tasks into a complete, iterative refinement loop, where mutually enhancing shape and topology prediction is achieved, thereby improving token prediction accuracy. 2) An improved Hourglass architecture designed to process and leverage multi-level mesh information effectively. This architecture integrates Multi-level RoPE and an enhanced transformer block to capture crucial hierarchical geometric context, supporting highly accurate token prediction across long and spatially complex sequences. 3) A novel connection loss that directly introduces topology priors during training. This loss enforces hard topology constraints by ensuring consistency for tokens sharing the same vertex, thereby significantly enhancing the model’s ability to maintain topological integrity. Collectively, these three components ensure that TSSR fully utilizes the rich information available in parallel token generation. This strategic integration significantly improves token prediction accuracy, thereby addressing the loss of sequential causal dependencies and rendering parallel mesh generation highly effective for artist-style meshes. Experimental results on various open datasets demonstrate TSSR’s successful generation of artist-style meshes via parallel token generation at high resolutions (i.e., $1024^{3}$) up to 10K faces.
@@
  • By carefully designing, TSSR enables full leverage of the rich information after seeing all tokens of models in token parallel generation.
@@ 2 Related Work
Shape-then-mesh generation. These works focus on first generating the shape via a continuous volumetric field, where SDF is the most common choice. Then, SDF can be converted to the mesh via the Marching Cube [15]. For example, TRELLIS [25] and Hunyuan 2.5 [11] first leverage the VAE to reconstruct the mesh via SDF. Then, they leverage the diffusion model to generate the VAE latent. The advantage of this is that the explicitly 3D information can be introduced into the VAE. Diffusion itself can be more robust to the 3D shape. The difference is that TRELLIS introduces the explicit shape constraint into VAE training. Hunyuan proposed the improved version to extract GT SDF. Besides enhancing VAE, Wonder3D [14] introduces a cross-domain diffusion model framework to further enhance the diffusion process.
@@
Except for the ARs, there is some exploration for token parallel generation via diffusion. For example, Meshcraft [9] and Polydiff [1] proposed directly mapping the mesh into a continuous latent space via VAE and then implementing diffusion models on the latent space. However, the compression loss of VAE will inevitably lead to the degradation again.
Discrete Diffusion Models. DDMs [7, 19] are another type of diffusion model compared to the continuous state-based diffusion models [13]. Previous works have begun to implement DDMs on various tasks, such as LLM [21] and image generation [23], demonstrating the potential of DDMs. In this work, we propose TSSR, a token parallel generation paradigm based on the DDM. TSSR directly generates the artist-like method without any conversion, such as SDF and VAE, thereby maintaining the high quality.
@@ 4 Method
In this section, we demonstrate how TSSR effectively leverages the global shape and intricate topology available in parallel token generation. This is achieved through three integrated core contributions: 1) a unique decoupling training and hybrid inference, 2) an improved Hourglass architecture, and 3) the proposed connection loss.
@@
The iterative rule involves cycling through the topology sculptor $\theta_{mask}$ and $\theta_{uniform}$, as well as a classifier $\phi$, to decide which tokens to keep. We define $x_{t,mask}$ and $x_{t,uniform}$ to represent the mask-based noise corrupted and uniform-based noise corrupted $x_{t}$, respectively. Then, $x_{t}$, $t<1$ , always have mask noise and thus $x_{t,massk}=x_{t}$. We do the mask-based denoising process to calculate $p_{1|t}(x_{t,mask},t,c,\theta_{mask})$ and get its related clean state prediction $x_{1|t,mask}$. $x_{1|t,mask}$ will inevitably have some errors in token results due to the bias of $\theta_{mask}$, but keep some local intricate topology based on our motivation. Thus, it can be naturally regarded as adding a special type of uniform-based noise caused by model bias. This allows TSSR to use the shape refiner to correct these error tokens again and recover the shape, where we have $x_{t,uniform}=x_{1|t,mask}$. We then do uniform-based denoising process to calculate $p_{1|t}(x_{t,uniform},t,c,\theta_{uniform})$ and get its related clean state prediction $x_{1|t,uniform}$.
Now, the final step is to decide how many tokens should be kept and transfer the rest to [MASK] to get $x_{t+\Delta t}$. We leverage $\phi$ to calculate:
@@
Then, for the token position, we proposed Multi-level RoPE. The non-AR nature of DDMs allows every token to attend to every other token, creating an opportunity to inject richer structural information than is possible in AR models. Standard positional embeddings, such as 1D RoPE, treat the mesh sequence as a flat list, ignoring its inherent hierarchical structure (coordinates within vertices, vertices within faces). To capitalize on DDM’s full attention scope, we propose a novel multi-level RoPE that explicitly encodes the mesh’s hierarchical structure, shown in Fig. 3 (c). Instead of a single position index, each token is assigned a 3D-like positional tuple: $(\text{face\_index},\text{vertex\_index},\text{coordinate\_index})$. This allows the model’s attention mechanism to learn relationships not just between adjacent tokens in the sequence, but also between vertices within the same face, or across different faces, providing a powerful inductive bias for learning topology.
In the end, we show a trick to unify three models into one with a light-aware classifier to leverage the latent from the transformer. We first added a new condition called the conditional task flag, noted as $c_{flag}$. The flag indicates the current task being performed (e.g., $c_{flag}=0$ for mask-based denoising, $c_{flag}=1$ for uniform-based denoising. This approach promotes parameter efficiency, allowing the model to learn shared representations of mesh structure that benefit both topology sculpting and shape refining tasks. This enables us to use the one model $\theta$ to cover the ability of $\theta_{mask}$ and $\theta_{uniform}$. Then, we integrate a light-aware classifier $\phi$ directly into the architecture. A lightweight classification head is attached to the final layer of the Hourglass transformer. This head takes the model’s final token representations as input and predicts the probability of each token being part of a topologically incorrect region. This shared representation ensures that the features learned for generation are effectively reused for error detection.
Connection loss. After enabling the token parallel generation, we find that topology can be represented by correctly repeating the token among shared vertices. Since each token represents a discrete XYZ token, shared vertices will have the same token. Therefore, we leverage this to propose the connection loss as a hard constraint to improve the accuracy of token prediction further.
@@
We summarize our hybrid inference and decoupling training via improved Hourglass in Algorithm 1 and Algorithm 2. In this way, TSSR successfully maintains accurate token prediction via a parallel token generation manner. Our design enables TSSR to generate meshes with high resolution and up to 10,000 faces.
@@
Algorithm 2 Decoupling training
1:Input: Epochs
2:Output: $\theta$$\phi$
3:Initialize $\theta$ and $\phi$
4:for step in $0,...,\text{Epochs}]$ do
5:  Sample $\hat{x}_{1}$ and $c$ from Dataset.
6:  $t\sim[0,1000]$.
7:  Calculate $x_{t,mask}$ via Eq. 2.
8:  Calculate $p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
9:  Calculate $\mathcal{L}_{mask}$ via Eq. 5.
10:  $x_{t,uniform}\sim p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
11:  Calculate $p_{1|t}(x_{t,uniform},t,c,c_{flag}$
12:  Calculate $\mathcal{L}_{phi}$ via Eq. 7.
13:  Calculate $L_{connection}$ via Eq. 12.
14:  $\theta,\phi\leftarrow\nabla_{\theta,\phi}\mathcal{L}$
15:end for
16:Return: $\theta$$\phi$
@@ 5.1 Implementation
Setting. We train a 2B model from scratch. Training is conducted over 20 days on a cluster of 32 NVIDIA A100 40GB GPUs, initially focusing on meshes with face counts in the range $[0,7000]$ to obtain an initial checkpoint. Then, it scales to $[7000,10000]$ face counts trained on the 7 days using a cluster of 16 NVIDIA H20 96GB GPUs. We use a linear learning rate scheduler with a warm-up strategy and the AdamW optimizer [16], increasing from $0,1\times 10^{-4}]$ over the first 100 steps and maintaining $1\times 10^{-4}$ thereafter. We set $T=200$ during inference.

### v2

@@ Abstract
In this paper, we introduce Topology Sculptor, Shape Refiner (TSSR), a novel method for generating high-quality, artist-style 3D meshes based on Discrete Diffusion Models (DDMs). Our primary motivation for TSSR is to achieve highly accurate token prediction while enabling parallel generation, a significant advantage over sequential autoregressive methods. By allowing TSSR to ”see” all mesh tokens concurrently, we unlock a new level of efficiency and control. We leverage this parallel generation capability through three key innovations: 1) Decoupled Training and Hybrid Inference, which distinctly separates the DDM-based generation into a topology sculpting stage and a subsequent shape refinement stage. This strategic decoupling enables TSSR to capture both intricate local topology and overarching global shape effectively. 2) An Improved Hourglass Architecture, featuring bidirectional attention enriched by face-vertex-sequence level Rotational Positional Embeddings (RoPE), thereby capturing richer contextual information across the mesh structure. 3) A novel Connection Loss, which acts as a topological constraint to further enhance the realism and fidelity of the generated meshes. Extensive experiments on complex datasets demonstrate that TSSR generates high-quality 3D artist-style meshes, capable of achieving up to 10,000 faces at a remarkable spatial resolution of $1024^{3}$. The code will be released at: https://github.com/psky1111/Tencent-TSSR.
@@ 1 Introduction
Mainstream 3D mesh generation methods can be broadly categorized into two approaches: 1) Shape-then-mesh generation and 2) direct mesh generation. Shape-then-mesh generations [28] start with representing a shape via continuous volumetric fields, such as Neural Radiance Fields (NeRF) [17] and Signed Distance Fields (SDF) [10]. Then, a diffusion model is used to generate such a representation and convert it to real meshes via Marching Cubes [15]. Benefits from parallel token generation in diffusion models, but shape-then-mesh generation is efficient, but suffers from inevitable denegation after the conversion process. Direct mesh generation methods aim to predict mesh vertices and faces directly. Mainstream methods, such as Meshtron [8] and DeepMesh [27], employ the Auto-Regressive (AR) approach, which first quantizes the mesh coordinates into a discrete space (V discrete bins for each coordinate, where V is also referred to as the spatial resolution in this context). Then it implements the specific tokenizer method to generate the mesh token by token. Benefits from the discrete space, ARs can generate high-quality meshes, but struggle with efficiency.
@@
In this paper, we propose the Topology Sculptor, Shape Refiner (TSSR), a novel paradigm explicitly designed for direct mesh generation built upon DDMs. TSSR addresses the challenge of maintaining accurate token prediction in parallel generation for meshes. TSSR leverages DDMs’ inherent ability to generate tokens in parallel, enabling the model to view all tokens simultaneously and to utilize bidirectional attention to capture comprehensive global shape and intricate topological information. TSSR realizes accurate parallel mesh generation through three integrated core contributions:
1) A unique decoupling training and hybrid inference paradigm that strategically divides mesh generation into specialized sub-tasks for shape (shape refiner) and topology (topology sculptor). This strategy acknowledges that while parallel generation yields rich, comprehensive information (encompassing global shape and intricate topology), effectively processing this intertwined complexity for accurate prediction is challenging. Decoupling training enables separating these aspects into distinct sub-tasks, thereby reducing the learning burden on the model and allowing each sub-task to focus predominantly on either shape or topology. Then, the hybrid inference process unifies these subtasks into a complete, iterative refinement loop, enabling mutually enhancing shape and topology prediction and thereby improving token prediction accuracy. 2) An improved Hourglass architecture designed to process and leverage multi-level mesh information effectively. This architecture integrates Multi-level RoPE and an enhanced transformer block to capture crucial hierarchical geometric context, enabling highly accurate token prediction across long, spatially complex sequences. 3) A novel connection loss that directly introduces topology priors during training. This loss enforces hard topology constraints by ensuring token consistency across vertices, thereby significantly enhancing the model’s ability to maintain topological integrity. Collectively, these three components ensure that TSSR fully utilizes the rich information available in parallel token generation. This strategic integration significantly improves token prediction accuracy, thereby addressing the loss of sequential causal dependencies and rendering parallel mesh generation highly effective for artist-style meshes. Experimental results on various open datasets demonstrate TSSR’s ability to generate artist-style meshes via parallel token generation at high resolutions (i.e., $1024^{3}$) and up to 10K faces.
@@
  • By carefully designing, TSSR enables full leverage of the rich information after ”seeing” all tokens in token-parallel generation.
@@ 2 Related Work
Shape-then-mesh generation. These works focus on first generating the shape via a continuous volumetric field, where SDF is the most common choice. Then, SDF can be converted to a mesh using the Marching Cubes [15]. For example, TRELLIS [25] and Hunyuan 2.5 [11] first leverage the VAE to reconstruct the mesh via SDF. Then, they leverage the diffusion model to generate the VAE latent. The advantage is that the explicit 3D information can be incorporated into the VAE. Diffusion itself can be more robust to the 3D shape. The difference is that TRELLIS introduces the explicit shape constraint into VAE training. Hunyuan proposed an improved version for extracting GT SDF. Besides enhancing VAE, Wonder3D [14] introduces a cross-domain diffusion model framework to further enhance the diffusion process.
@@
Except for the ARs, there is some exploration of token-parallel generation via diffusion. For example, Meshcraft [9] and Polydiff [1] proposed directly mapping the mesh into a continuous latent space via VAE and then implementing diffusion models on the latent space. However, the compression loss in VAE will inevitably lead to further degradation.
Discrete Diffusion Models. DDMs [7, 19] are another type of diffusion model compared to the continuous state-based diffusion models [13]. Previous work has begun to implement DDMs across various tasks, such as LLMs [21] and image generation [23], demonstrating their potential. In this work, we propose TSSR, a token parallel generation paradigm based on the DDM. TSSR directly generates artist-like methods without any conversion, such as SDF and VAE, thereby maintaining high quality.
@@ 4 Method
Problem definition. Our target is to generate artist-like meshes given the point cloud condition $c$ via DDMs to enable parallel token generation. The overall pipeline contains inference and training. During training, we can access GT meshes and related $c$. We convert GT meshes into quantized, discrete token sequences $\hat{x}_{1}$ using the tokenizer. Then, we add noise into the $\hat{x}_{1}$ to get $x_{t}$. In the end, we train a model to predict the clean token sequence. During inference, we cannot access GT meshes. We will initialize with arbitrary token sequences composed of pure noise tokens. Then, we iteratively denoise the noise via a trained model. In the end, we obtain a final clean token sequence $x_{1}$ and a de-tokenizer to generate meshes.
In this paper, TSSR effectively leverages the global shape and intricate topology available in parallel token generation to propose novel training and inference via three integrated core contributions: 1) a unique decoupling training and hybrid inference, 2) an improved Hourglass architecture, and 3) the proposed connection loss.
@@
The iterative rule involves cycling through the topology sculptor $\theta_{mask}$ and $\theta_{uniform}$, as well as a classifier $\phi$, to decide which tokens to keep. We define $x_{t,mask}$ and $x_{t,uniform}$ to represent the mask-based noise corrupted and uniform-based noise corrupted $x_{t}$, respectively. Then, $x_{t}$, $t<1$ , always have mask noise and thus $x_{t,massk}=x_{t}$. We do the mask-based denoising process to calculate $p_{1|t}(x_{t,mask},t,c,\theta_{mask})$ and get its related clean state prediction $x_{1|t,mask}$. $x_{1|t,mask}$ will inevitably have some errors in token results due to the bias of $\theta_{mask}$, but keep some local intricate topology based on our motivation. Thus, it can naturally be regarded as adding a special type of uniform-based noise due to model bias. This allows TSSR to use the shape refiner to correct these error tokens again and recover the shape, where we have $x_{t,uniform}=x_{1|t,mask}$. We then do uniform-based denoising process to calculate $p_{1|t}(x_{t,uniform},t,c,\theta_{uniform})$ and get its related clean state prediction $x_{1|t,uniform}$.
Now, the final step is to decide how many tokens to keep and transfer the rest to [MASK] to obtain $x_{t+\Delta t}$. We leverage $\phi$ to calculate:
@@
Then, for the token position, we proposed Multi-level RoPE. The non-AR nature of DDMs allows every token to attend to every other token, creating an opportunity to inject richer structural information than is possible in AR models. Standard positional embeddings, such as 1D RoPE, treat the mesh sequence as a flat list, ignoring its inherent hierarchical structure (coordinates within vertices, vertices within faces). To capitalize on DDM’s full attention scope, we propose a novel multi-level RoPE that explicitly encodes the mesh’s hierarchical structure, shown in Fig. 3 (c). Instead of a single position index, each token is assigned a 3D-like positional tuple: $(\text{face\_index},\text{vertex\_index},\text{coordinate\_index})$. This allows the model’s attention mechanism to learn relationships not only between adjacent tokens in the sequence, but also between vertices within the same face and across different faces, providing a powerful inductive bias for learning topology.
In the end, we show a trick to unify three models into one with a light-aware classifier to leverage the latent from the transformer. We first added a new condition, the conditional task flag ($c_{flag}$). The flag indicates the current task being performed (e.g., $c_{flag}=0$ for mask-based denoising, $c_{flag}=1$ for uniform-based denoising. This approach promotes parameter efficiency, allowing the model to learn shared representations of mesh structure that benefit both topology sculpting and shape refining tasks. This enables us to use the single model $\theta$ to capture the abilities of $\theta_{mask}$ and $\theta_{uniform}$. Then, we integrate a light-aware classifier $\phi$ directly into the architecture. A lightweight classification head is attached to the final layer of the Hourglass transformer. This head takes the model’s final token representations as input and predicts the probability that each token belongs to a topologically incorrect region. This shared representation ensures that the features learned for generation are effectively reused for error detection.
Connection loss. After enabling the token parallel generation, we find that topology can be represented by correctly repeating the token among shared vertices. Since each token represents a discrete XYZ token, shared vertices will have the same token. Therefore, we leverage this to propose using connection loss as a hard constraint to further improve token prediction accuracy.
@@
We summarize our hybrid inference and decoupled training via an improved Hourglass in Algorithms 2 and 1. In this way, TSSR successfully maintains accurate token prediction via parallel token generation. Our design enables TSSR to generate high-resolution meshes with up to 10,000 faces.
Algorithm 1 Decoupling training
1:Input: Epochs
2:Output: $\theta$$\phi$
3:Initialize $\theta$ and $\phi$
4:for step in $0,...,\text{Epochs}]$ do
5:  Sample $\hat{x}_{1}$ and $c$ from Dataset.
6:  $t\sim[0,1000]$.
7:  Calculate $x_{t,mask}$ via Eq. 2.
8:  Calculate $p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
9:  Calculate $\mathcal{L}_{mask}$ via Eq. 5.
10:  $x_{t,uniform}\sim p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
11:  Calculate $p_{1|t}(x_{t,uniform},t,c,c_{flag}$
12:  Calculate $\mathcal{L}_{phi}$ via Eq. 7.
13:  Calculate $L_{connection}$ via Eq. 12.
14:  $\theta,\phi\leftarrow\nabla_{\theta,\phi}\mathcal{L}$
15:end for
16:Return: $\theta$$\phi$
@@ 5.1 Implementation
Setting. We train a 2B model from scratch. Training is conducted over 20 days on a cluster of 32 GPUs with 40GB MEM, initially focusing on meshes with face counts in the range $[0,7000]$ to obtain an initial checkpoint. Then, it scales to $[7000,10000]$ face counts trained on the 7 days using a cluster of 16 GPUs with 96GB MEM. We use a linear learning rate scheduler with a warm-up strategy and the AdamW optimizer [16], increasing from $0,1\times 10^{-4}]$ over the first 100 steps and maintaining $1\times 10^{-4}$ thereafter. We set $T=200$ during inference.